This page represents notes on work to maintain monitors for the Broadcast Television Engineering Technology (BTV) program studio.
The report for the time I spent in the monitors group, including time as guru, is here, in the following formats:
Working on the monitor trio (Sony PVM-5300) for the video server workstation in the control room. Hooking in a tri-color (RGB) LED on front of each monitor to replace existing neon tally light. Using electrical tape (a real kludge) to hold the LEDs in place, cutting out a metal bracket that was holding a plastic mounting jig for the neon lights being replaced (continuing a decision made by the previous team). The plastic lenses in front of the LEDs have to be drilled out to make room for the LEDs since we are mounting them pointing towards the front of the units (this is how the previous team decided to do this). Drilled holes in access plate on back, hooked up three jacks (one for each color) to enable them.
The hot lead is hooked to the hot wire on the LED through a 1 megohm (later found to be about 800 ohms) resistor on Q board point R-227, on the side of the resistor R-227 that is closer to J5 and point 7. The ground leads on the jacks ganged together, wired to D board at point 7.
The LED pinout, in order, is red, power, green, blue. The power lead is the longest of the four leads, so you'd better know which is which before trimming the leads. Fortunately, the power lead also looks directly connected to large silver metal areas within the clear-plastic-surrounded area of the LED, so you can still tell what lead goes to what even after you trim them.
The LED requires 1.2 volts at 20 mA.
The jacks mounted on the back of the monitors are 3-conductor RCA (stereo - tip, ring, sleeve). We are using 3-conductor instead of 2-conductor because they were the same price, according to Steve. We only need 2-conductor RCA. Use a 1/4 inch drill bit for making holes for the 3-conductor jacks. When drilling holes in an access plate, place the holes farther down to make room to avoid cutting the plastic insulator board behind the access plate.
Wired LED into monitor.
Dan completed the monitor LED wiring on the second monitor. Work on the first monitor had already been completed by the previous team.
The LED doesn't light. Voltage at Q board point R-227 is 18V. On other side of 1 megohm resistor, voltage at LED hot leg is about 16 V. Turns out we got the wrong information, likely from a misunderstanding. We did the Ohm's law math and decided to use an 820 ohm 1/2 watt resistor to limit current and drop voltage. We actually used a larger wattage (probably 1 watt) on the second monitor just because we figured the resistor would need to dissipate 1/3 watt and thus we could have a better safety margin.
This change in resistor resulted in the LED lighting up as expected when inserting the shorted test plug into the jacks.
Dan gets transferred out to the intercom group. Brian Hughes is transferred into the monitors group as the apprentice. I am now the guru after one week on the monitors group.
Brian realized the Molex power connectors are wired backward with respect to gender. The female connector should have power when disconnected because the female has recessed metal while the male has exposed metal pins; this is standard safety procedure as evidenced by the computer power cord standard, and the fact that electrical power receptacles are jacks not plugs! We decided to change the gender of the connectors to correct this safety issue.
To keep heat shrink tubing from shring at one end where can't just add more tubing on that side (maybe because it butts up against a much bigger diameter object), heat the end of the heat shrink tubing that is furthest away from the big object.
Monitor power wires have no grounding connection when operated disconnected from the trio unit chassis. The chassis was connecting each monitor's grounding system with the green grounding wire in the power cord. The cheater cord was three wire but the monitor's Molex was only two wire. We decided to add green grounding wires to each monitor between the middle pin of the Molex connector and a suitable location on the respective monitor's chassis.
We discovered the cheater cord was wired with female Molex pins in a plug where male pins should be. We decided to correct that as part of correcting the Molex connector gender described earlier.
We completed rewiring two monitors.
Now we are working on the middle monitor, and adding a Molex connector for the left-side monitor at the chassis.
There is a 3.7 megohm (actual measured value 3.0 megohm) resistor mounted on the terminal strip handling power wires, between the white grounded wires and the chassis and green grounding wire. We wondered why it was there. Maybe to keep transformer primaries ungrounded. Maybe it is to discharge static electricity; Steve said one usually sees a capacitor or a capacitor plus resistor at this point.
Completed wiring of female Molex connector from power to monitor 3.
Diagrammed power terminal strip wiring from chassis back and top. I'm not sure if I will add this diagram to this log on my web site!
Installed LEDs, fastened bottom plates and covers, finished main power Molex off terminal strip, assembled monitor trio into chassis.
Re-orienting transversely-mounted LEDs to axial orientation (LEDs pointed toward the monitor front, the same orientation as is the CRT) for increased brightness.
This completes the job of rewiring the monitor trio to use the three-color LEDs for tally lights.
Aligned the left-side monitor (#1).
Aligned the middle monitor (#2). Instructions follow.
Here is how to align one of the monitors in the trio:
Here is how to perform white balancing of monitors using the color analyzer. It is so much nicer to use the color analyzer than it was attempting to white balance the monitors in the monitor trio using the service manual instructions because the sensor takes much of the subjectivity out of the measurement.
Work on the monitor for the Mac running Final Cut Pro. Sony BVM-8044QD. Doing this just because I'm spending time using that monitor editing the outreach video. We will do the bridge monitors next. It's easier to adjust a much more modern monitor, and to do it on the bench rather than on a ladder at the bridge, just so we get some more practice before we move to the bridge monitors.
Brought the monitor to the lab and powered it up first thing so it could warm up.
Powered up the monitor trio.
Installed monitor trio. Looks better overall. Mounting method seems weak: the plywood underneath the control room bridge counter is splitting from the wood screws we used.
Worked on the BVM-8044QD.
Saw some blue in some corners after degaussing. Probably was there before; just a problem with this monitor.
Tried to adjust purity. Purity rings, static convergence magnets all grouped together; hard to adjust one set without moving the other set. This adjustment didn't seem to do much. We just did the red color. You don't need to slide the deflection yoke back much to get the blue/green/red vertical stripes in the monitor. The instructions said to switch the pattern generator to green. We plugged in the flat field to G of the RGB ports and set the switch on the back to RGB (not Y/R-Y/B-Y). We moved the round flat magnet in the lower right corner of the CRT a little to improve the lower right corner purity, which had been a bit blue.
Tried to adjust convergence and geometry. The manual specifies to use the registration chart with the circles in the corners, the one with the Indian head. We used the little Sony pattern generator to give us this chart, except it had a dragon in the place of the Indian head, which presented a huge problem (just kidding). The vertical and horizontal size and centering had the greatest effect, since some of the frame was being cut off before.
To prepare for the convergence adjustment, did D-board adjustments as specified in the manual.
Convergence. H Stat VR, V stat magnet. HMC and VMC looked OK. Dynamic adjustment looks like requires spacers we don't have.
Performed focus adjustment.
Performed white balance.
Measured RGB cutoff for 103V. Green was about 93V. The meter seemed to load down the circuits so we didn't think we could perform an adjustment of this with confidence.
Reassembled the monitor and reinstalled it in its place within the Mac Final Cut Pro editing system.